OpenStudy (anonymous):

What are the prime factors of 18? 2, 9 3, 6 2, 3, 6, 9 2, 3, 3

OpenStudy (anonymous):

@laishaleecolon

OpenStudy (tkhunny):

One at a time. Keep sotting them until there are none left to spot. 18 What comes to mind first? Maybe, "Hey, that's even. There's a 2." 18 = 2*9 Now what? "9", that's jsut 3*3 18 = 2*3*3 Now what? Hmmm... Those are all prime. I think I'm done.

OpenStudy (anonymous):

@lexiamee

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

i am confused

hero (hero):

^What @tkhunny said

OpenStudy (anonymous):

i dont understand what she said is it 2,9

hero (hero):

In order to understand what she said you have to know your multiplication tables and understand what a prime number is.

OpenStudy (anonymous):

i know the multiplication table

OpenStudy (anonymous):

all the numbers are prime numbers to 18 i think it 2,3,69 because that has all the answers

OpenStudy (anonymous):

i ment 6,9

OpenStudy (anonymous):

the answer is 2,3,6,9 right

OpenStudy (anonymous):

THE ANSWER IS 2,3,3

hero (hero):

@anas2000, you need to study prime numbers. If you don't you will be lost forever

OpenStudy (anonymous):

the reason its not 2,3,6,9 because that would be the answer 324 not 18, do you get it now, you have to multiply those numbers to see if it comes out to be the number 18.

OpenStudy (anonymous):

2,3,3 come out to be the answer 18. =)

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

your welcome anytime=)

OpenStudy (anonymous):

please start practicing more the prime numbers, ok.

OpenStudy (anonymous):

the prime numbers of all the numbers

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

=)